OpenStudy (anonymous):

Consider the unbalanced equation for the decomposition of silver nitrate: AgNO3(s) --->Ag(s)+NO2(g)+O2(g) 1. Write the balanced equation for the reaction. 2. When 54.0g of AgNO3 is heated, 4.60g of O2 is formed. How many grams of Ag is also formed?

OpenStudy (anonymous):

2AgNO3(s) --->2Ag(s) + 2NO2(g) + O2(g) Moles of AgNO3: 2 * 4.6/32 = .288 mol Mass of Ag: .288 * 108 = 31.05 g

OpenStudy (anonymous):

can you help me with this please. What is the percent yield of Ag for the reaction?

OpenStudy (anonymous):

Theoretical mole of AgNO3: 54/ 170 = .318 mol Theoretical mass of Ag: .318 * 108 = 34.31g Percent yield: 31.05 * 100 / 34.31 = 90.51%